# SuperMat: Physically Consistent PBR Material Estimation at Interactive Rates

### v3

@@ 7 Additional Visualizations
Figure 14: Additional results on real-world objects. The objects are extracted from captured images by SAM [19].
Albedo Metallic Roughness Relighting
L1 26.3745 0.9159 0.0987 21.8087 0.8733 0.2113 22.4262 0.9013 0.1308 25.8089 0.9300 0.0765
SSI 23.8053 0.9073 0.1067 21.6018 0.8378 0.2097 22.8090 0.9063 0.1261 23.8945 0.9242 0.0813
Perceptual 27.7214 0.9252 0.0771 23.1332 0.8799 0.1584 24.7161 0.9154 0.0926 28.0626 0.9424 0.0548
@@
8 Additional Experiments
We show the detailed quantitative comparison in Tables 345 and  6. Additionally, without re-render loss, we experiment with three types of loss functions in the end-to-end framework: L1 loss, shift and scale invariant (SSI) loss [31], and perceptual loss. The performance of models trained with these loss designs is evaluated on the same test dataset, and the results are shown in Tab. 2. Among these losses, the perceptual loss, which is ultimately adopted by SuperMat, demonstrates the best performance.
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
Derender3D 0.0094 22.0689 0.8529 0.1754 0.0104 22.4523 0.8919 0.1489
IIR 0.0130 21.3350 0.8705 0.1587 0.0202 20.5980 0.8654 0.1672
IID 0.0081 23.3557 0.8916 0.1212 0.0122 22.3877 0.8845 0.1395
@@ 8 Additional Experiments
IntrinsicAnything 0.0121 21.3502 0.8743 0.1481 0.0219 20.2352 0.8700 0.1682
StableMaterial 0.0101 23.8964 0.9038 0.0985 0.0143 23.9020 0.9061 0.1079
Ours w/o e2e 0.0058 24.9261 0.8997 0.0986 0.0102 24.3049 0.8935 0.1219
Ours w/o re-render 0.0023 27.7214 0.9252 0.0771 0.0043 26.7084 0.9185 0.0953
Ours 0.0020 28.1156 0.9284 0.0753 0.0044 26.8911 0.9195 0.0938
@@
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0471 16.1588 0.8401 0.1997 0.0515 17.6046 0.7835 0.2732
IID 0.0343 17.3157 0.8430 0.1954 0.0431 17.4303 0.7742 0.2866
@@
StableMaterial 0.0450 16.6922 0.8487 0.2034 0.0515 20.7938 0.8156 0.2379
Ours w/o e2e 0.0353 18.7623 0.8696 0.1879 0.0470 20.9061 0.8455 0.2764
Ours w/o re-render 0.0134 23.1332 0.8799 0.1584 0.0330 22.7159 0.8150 0.2352
Ours 0.0125 23.5208 0.8845 0.1573 0.0291 23.3987 0.8351 0.2302
@@
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0195 20.7515 0.8877 0.1367 0.0334 19.1632 0.8621 0.1790
IID 0.0176 21.5627 0.8917 0.1174 0.0295 19.9655 0.8718 0.1586
@@
StableMaterial 0.0133 21.9898 0.9096 0.1024 0.0251 20.9379 0.8962 0.1312
Ours w/o e2e 0.0161 21.1543 0.8858 0.1148 0.0278 20.4794 0.8679 0.1633
Ours w/o re-render 0.0070 24.7161 0.9154 0.0926 0.0221 21.9006 0.8933 0.1317
Ours 0.0070 24.7642 0.9144 0.0915 0.0214 22.2674 0.8935 0.1313
@@
Objaverse BlenderVault
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0124 22.5646 0.9021 0.1116 0.0296 20.4861 0.8782 0.1441
IID 0.0085 24.3354 0.9144 0.0903 0.0160 22.3058 0.9009 0.1142
@@
StableMaterial 0.0115 23.5445 0.9121 0.0814 0.0203 22.2230 0.9022 0.1013
Ours w/o e2e 0.0072 25.5109 0.9219 0.0739 0.0171 23.1506 0.9031 0.1059
Ours w/o re-render 0.0041 28.0626 0.9424 0.0548 0.0123 25.5175 0.9259 0.0826
Ours 0.0035 28.4900 0.9465 0.0510 0.0106 25.7484 0.9266 0.0770

### v4

@@ 10 Limitations
Figure 13: Additional albedo comparison from the DTC test dataset.
@@
Figure 16: Additional metallic and roughness comparison from the DTC test dataset. The metallic maps are shown on the left side (M), while the roughness maps are shown on the right side (R).
@@
Objaverse BlenderVault DTC
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
Derender3D 0.0087 22.16 0.8532 0.1816 0.0104 22.45 0.8919 0.1489 0.0246 18.30 0.8366 0.2157
IIR 0.0100 22.99 0.8741 0.1396 0.0150 22.32 0.8709 0.1616 0.0165 20.50 0.8677 0.1644
IID 0.0088 23.15 0.8847 0.1276 0.0122 22.39 0.8845 0.1395 0.0101 22.15 0.9031 0.1175
@@
IntrinsicAnything 0.0125 21.36 0.8701 0.1550 0.0219 20.24 0.8700 0.1682 0.0122 21.23 0.8963 0.1328
StableMaterial 0.0110 23.78 0.8989 0.1064 0.0143 23.90 0.9061 0.1079 0.0129 22.64 0.9008 0.1086
StableMaterialMV 0.0074 24.41 0.8974 0.0970 0.0089 25.13 0.9071 0.1014 0.0067 24.15 0.9229 0.0855
SuperMat w/o e2e 0.0057 25.11 0.8973 0.1027 0.0102 24.30 0.8935 0.1219 0.0107 23.36 0.8985 0.1038
SuperMat w/o re-render 0.0029 27.01 0.9151 0.0949 0.0047 26.12 0.9152 0.1045 0.0031 26.96 0.9405 0.0725
SuperMat 0.0024 27.66 0.9209 0.0865 0.0044 26.63 0.9171 0.0997 0.0018 28.74 0.9490 0.0597
SuperMatMV 0.0022 28.04 0.9241 0.0832 0.0039 26.75 0.9183 0.1026 0.0023 27.89 0.9443 0.0658
@@
Objaverse BlenderVault DTC
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0462 16.87 0.8399 0.2230 0.0472 17.11 0.7688 0.2900 0.0324 19.86 0.6771 0.3604
IID 0.0354 17.31 0.8307 0.2108 0.0431 17.43 0.7742 0.2866 0.0347 17.41 0.6797 0.3895
@@
StableMaterial 0.0493 16.83 0.8398 0.2164 0.0515 20.79 0.8156 0.2379 0.0567 23.26 0.7938 0.2833
StableMaterialMV 0.0452 16.96 0.8411 0.2114 0.0511 18.74 0.8150 0.2590 0.0777 16.15 0.7096 0.3783
SuperMat w/o e2e 0.0373 19.40 0.8672 0.2021 0.0470 20.91 0.8455 0.2764 0.0496 22.05 0.8204 0.3694
SuperMat w/o re-render 0.0135 22.87 0.8669 0.1760 0.0311 22.52 0.8040 0.2343 0.0087 28.23 0.8067 0.2687
SuperMat 0.0109 23.78 0.8785 0.1695 0.0344 23.02 0.8335 0.2328 0.0068 29.65 0.8936 0.2641
SuperMatMV 0.0069 25.38 0.8919 0.1619 0.0324 23.17 0.8437 0.2311 0.0059 29.78 0.8892 0.2656
@@
Objaverse BlenderVault DTC
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0257 20.67 0.8902 0.1242 0.0438 18.58 0.8503 0.1754 0.0320 19.93 0.8512 0.1793
IID 0.0204 21.32 0.8863 0.1266 0.0295 19.97 0.8718 0.1586 0.0168 21.51 0.8695 0.1576
@@
StableMaterial 0.0131 21.97 0.9065 0.1064 0.0251 20.94 0.8962 0.1312 0.0256 20.13 0.8612 0.1604
StableMaterialMV 0.0142 21.26 0.8949 0.1081 0.0253 20.49 0.8915 0.1284 0.0233 20.83 0.8651 0.1493
SuperMat w/o e2e 0.0170 21.36 0.8832 0.1221 0.0278 20.48 0.8679 0.1633 0.0246 20.60 0.8352 0.1907
SuperMat w/o re-render 0.0081 24.15 0.9145 0.1156 0.0211 21.73 0.8942 0.1429 0.0086 24.69 0.8961 0.1502
SuperMat 0.0074 24.59 0.9154 0.1114 0.0201 22.39 0.8972 0.1386 0.0053 25.78 0.9046 0.1342
SuperMatMV 0.0074 24.96 0.9165 0.1070 0.0145 23.22 0.8998 0.1297 0.0049 26.34 0.9064 0.1257
@@
Objaverse BlenderVault DTC
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
IIR 0.0132 23.19 0.8991 0.1060 0.0290 20.96 0.8765 0.1441 0.0245 18.80 0.8897 0.0890
IID 0.0105 23.83 0.9057 0.0971 0.0160 22.31 0.9009 0.1142 0.0149 21.21 0.9066 0.1137
@@
StableMaterial 0.0107 24.08 0.9108 0.0851 0.0203 22.22 0.9022 0.1013 0.0183 21.38 0.8999 0.1074
StableMaterialMV 0.0091 24.02 0.9114 0.0816 0.0154 22.69 0.9019 0.1018 0.0134 22.70 0.9137 0.0966
SuperMat w/o e2e 0.0079 25.81 0.9192 0.0776 0.0171 23.15 0.9031 0.1059 0.0158 22.75 0.9032 0.1057
SuperMat w/o re-render 0.0042 27.25 0.9374 0.0650 0.0109 24.94 0.9270 0.0855 0.0050 27.05 0.9490 0.0600
SuperMat 0.0041 28.01 0.9406 0.0566 0.0101 25.50 0.9300 0.0811 0.0031 29.47 0.9590 0.0460
SuperMatMV 0.0032 28.51 0.9437 0.0566 0.0118 25.41 0.9289 0.0815 0.0041 29.01 0.9567 0.0502